### v4

# Quantum metrological capability as a probe for quantum phase transition

Abstract
The comprehension of quantum phase transitions (QPTs) is considered as a critical foothold in the field of many-body physics. Developing protocols to effectively identify and understand QPTs thus represents a key but challenging task for present quantum simulation experiments. Here, we establish a dynamical quench-interferometric framework to probe a zero-temperature QPT, which utilizes the evolved state by quenching the QPT Hamiltonian as input of a unitary interferometer. The metrological capability quantified by the quantum Fisher information shows a unique peak in the vicinity of the quantum critical point, allowing us to probe the QPT without cooling the system to its ground state. We show that the probing can be implemented by extracting quantum fluctuations of the interferometric generator as well as parameter estimation uncertainty of the interferometric phase, and subsequently allows identifying the boundary of the phase diagram. Our results establish an important link between QPTs and quantum metrology, and enrich the toolbox of studying non-equilibrium many-body physics in current quantum simulators.
April 2, 2025
Introduction
Quantum phase transitions (QPTs) represent a vigorous subject strongly relevant in various fields, ranging from condensed matter to quantum field theory and cosmology [1, 2]. It is commonly considered an essential ingredient for our perception of a plethora of fascinating quantum phenomena in strongly correlated many-body physics such as quantum magnetization [3], spin liquid [4], topological order [5] and superconductor-insulator transition [6]. In the last few years, the booming development of AMO physics and quantum simulation technologies opens a pristine and versatile experimental avenue towards the study of QPTs through controllable realizations of a variety of theoretical models [7, 8, 9, 10, 11, 12, 13]. This also naturally inspires intensive attention of synergistically investigating QPTs in quantum simulators beyond the traditional Landau-Ginzburg paradigm, by resorting to the concepts and strategies developed in quantum information science [14, 15, 16, 17, 18].
Remarkably, a quantum-fidelity approach to QPTs has been put forward to unveil sharp variations of the ground state close to critical point from a universal information-theoretic picture [19, 20, 21, 15, 16]. However, the requirement of cooling a complex many-body system to nearly zero temperature might be challenging. More recently, frameworks based on non-equilibrium dynamical response after sudden quantum quenches interestingly demonstrate the possibility of establishing a versatile toolbox for locating and witnessing QPTs [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34], and are expected to become a routine practice. While being compelling, efficient extraction of the quantities that act as indicators of the QPTs in those previous frameworks, such as short-range correlations [22, 23], out-of-time-order correlators [24, 25], entanglement entropy [23] and even multi-site entanglement [28], typically requires the ability of implementing local measurements or other complex operations (e.g. time-reversed evolution) on a generic many-body system. Therefore, further strengthening and enriching the toolbox of probing QPTs with quantum information strategies still remains largely unexplored and deserves intense efforts.
In this Letter, we establish an efficient framework for probing QPTs by integrating nonequilibrium quench dynamics with the concept of interferometry in quantum metrology. Specifically, we consider the post-quenched state prepared by time evolution governed by a QPT Hamiltonian as input of a unitary interferometer. We stress that our interferometric framework is different from the typical scenarios of critical metrology, which focus on estimating the QPT’s control parameter that drives the phase transition [19, 35, 36, 37, 38, 39]. The metrological capability of the state, theoretically quantified by the quantum Fisher information, determines the ultimate precision of estimating the interferometric phase. As our main result, we find that the metrological capability interestingly shows a peak around the quantum critical point (QCP). We further demonstrate that this enhanced metrological capability can be effectively accessed through quantum fluctuation of the interferometric generator or the estimation uncertainty of the phase, which merely requires the measurement of certain collective spin operators by global detection of the QPT system. As a direct application, we obtain the boundary of the QPT phase diagram and show that it coincides well with the theoretical prediction of equilibrium QCPs. Our framework not only provides a new promising and experimentally practical avenue to investigate QPTs and the comprehension of non-equilibrium many-body quantum phenomena, but also reveals further deep connections between quantum metrology and quantum criticality [35, 36, 37, 38, 40, 41, 42, 39, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53].
Quench-interferometric framework for probing QPTs
At the heart of quantum metrology lies the quantum Fisher information (QFI)  [54, 55, 56], which can quantify the metrological capability of a probe state with respect to the unitary parameterization generated by a Hermitian operator $\hat{\mathcal{O}}$. In this work, we focus on the standard interferometric framework [57, 58, 59, 41, 60], where a local $N$-particle generator $\hat{\mathcal{O}}=\sum_{i=1}^{N}\hat{O}_{i}$ is considered. Given a pure probe state |$\ket{\Psi}$, the associated QFI can be easily computed or measured from quantum fluctuation of the generator [60, 56]
$F_{Q}[\ket{\Psi},\hat{\mathcal{O}}]=4\left[\langle\Psi|\hat{\mathcal{O}}^{2}|\Psi\rangle-\langle\Psi|\hat{\mathcal{O}}|\Psi\rangle^{2}\right].$ (1)
This form of QFI is elegantly linked to the connected correlation function, resulting in significant implications well beyond the scope of quantum metrology. Recently, in paradigmatic spin models of QPTs, it has been demonstrated that the QFI of the ground state can exhibit superextensive divergence or abrupt changes in the vicinity of a QCP [60, 41, 61]. Such fundamentally interesting behaviors could unveil a variety of fascinating phenomena such as the witness of critical multipartite entanglement [60], the redistribution of quantum noise among extensive observables [41] and the emergence of many-body nonlocality [61], which however is hindered by the requirement of ground state preparation.
Figure 1: Schematic graph of the quench-interferometric framework for probing QPTs. A many-body system of $N$ particles is prepared into a probe state via nonequilibrium quantum dynamics, achieved by suddenly quenching the QPT Hamiltonian $H_{\mathrm{QPT}}$. This probe state is then used to estimate an unknown phase $\vartheta$ encoded by a local generator $\hat{\mathcal{O}}=\sum_{i=1}^{N}\hat{O}_{i}$. The metrological capability of the probe state can be significantly enhanced when the quenched Hamiltonian approaches the quantum critical point, as evidenced by a much narrower probability density function (PDF) of the estimated value of $\vartheta$.
Here, we consider the QFI of a probe state prepared through nonequilibrium quantum dynamics by suddenly quenching the QPT Hamiltonian, thus circumventing the challenging preparation of a many-body system near its ground state. It is worth to remark that the time-dependent QFI evolution in quench dynamics might open new possibilities beyond the knowledge of quantum systems at equilibrium, enabling alternative pathways towards various intriguing phenomena such as entangling dynamics [62, 63, 64, 47, 49, 65], transient quantum coherence [66, 67] and the interplay between equilibrium and dynamical QCPs [22]. Explicitly, our study proposes a quench-interferometric framework that employs post-quenched states as input of interferometers, namely
$|\Psi_{t}(\vartheta)\rangle=e^{-i\vartheta\hat{\mathcal{O}}}e^{-iH_{\mathrm{QPT}}t}|\Psi_{0}\rangle.$ (2)
We firstly evolve the system to a probe state by a QPT Hamiltonian, i.e. $|\Psi_{t}\rangle=e^{-iH_{\mathrm{QPT}}t}|\Psi_{0}\rangle$, and then input it into a quantum interferometer with a phase unitarily encoded by $e^{-i\vartheta\hat{\mathcal{O}}}$, see Fig. 1. The ultimate measurement precision of the interferometric phase $\vartheta$ is determined by the QFI in $|\Psi_{t}(\vartheta)\rangle$, which is uniformly (namely, independent of $\vartheta$) given by $F_{Q}(t)=F_{Q}[|\Psi_{t}\rangle,\hat{\mathcal{O}}]$.
Below, we investigate identification of the QPT by analyzing QFI evolution of the post-quenched state in Eq. (2). As an illustrative example, we focus on the zero-temperature paramagnetic-to-ferromagnetic QPT embedded in the anisotropic next-nearest-neighbor Ising (ANNNI) chain, see Fig. 2(a), which is a representative model describing generic (quantum chaotic) many-body systems [68, 23]. The model Hamiltonian reads,
$H(\kappa,B)=-J\sum_{i=1}^{N}\sigma^{x}_{i}\sigma^{x}_{i+1}+\kappa\sum_{i=1}^{N}\sigma^{x}_{i}\sigma^{x}_{i+2}-B\sum_{i=1}^{N}\sigma^{z}_{i},$ (3)
where $\sigma^{x,z}_{i}$ are the Pauli operators at site $i$, and a periodic boundary condition is employed with $\sigma_{N+1}^{x}=\sigma_{1}^{x}$, $\sigma_{N+2}^{x}=\sigma_{2}^{x}$. We set $J=1$ in the following for simplicity. This model is generally nonintegrable [68, 69] except for the special scenario of $\kappa=0$. Second-order perturbation theory determines the paramagnetic-to-ferromagnetic transition line of the ANNNI chain as $1-2\kappa_{c}=B_{c}-\kappa_{c}B_{c}^{2}/(2-2\kappa_{c})$ [70]. In our framework, we initialize this spin chain in a fully polarized state $|\Psi_{0}\rangle=|{\uparrow}\rangle^{\otimes N}$ and let it evolve under the ANNNI Hamiltonian, i.e.
|$\ket{\Psi_{t}}=e^{-iH(\kappa,B)t}\ket{\Psi_{0}}.$ (4)
We consider the collective spin operator $\hat{\mathcal{O}}=\hat{S}_{x}=\sum_{i=1}^{N}\sigma_{i}^{x}/2$ as the phase-translation generator. Numerical calculations of $F_{Q}(t)=F_{Q}[\ket{\Psi_{t}},\hat{S}_{x}]$ for 20 spins are performed using the exact diagonalization (ED) method [71, 72]. In Fig. 2 (b-c), we demonstrate that the QFI evolution of |$\ket{\Psi_{t}}$ prepared by non-equilibrium quench dynamics in Eq. (4) can be employed to probe the QPT. It can be seen in Fig. 2 (b) that the QFI dynamics close to the QCP is radically enhanced in contrast to the scenarios away from the QCP. The averaged QFI is calculated over a time window of $T$ to quantify the overall difference across the critical point, i.e.
$\overline{F_{Q}}=(1/T)\int_{0}^{T}F_{Q}(t)\,\mathrm{d}t.$ (5)
Remarkably, $\overline{F_{Q}}$ displays a peak in the vicinity of the QCP as shown in Fig. 2 (c), which serves an effective indicator to locate the QPT. We note that the emergence of such a peak can be related to the spontaneous symmetry breaking [73]. Our numeric analysis also shows that these results are robust to a finite system size and open boundary condition [73].
Figure 2: Probing the zero-temperature QPT of the ANNNI model. (a) The ANNNI chain consists of nearest-neighboring and next-nearest-neighboring Ising interaction as well as a transverse biased field, denoted by the parameters $(J,\kappa,B)$ respectively. In our simulation we set $J=1$ and $B=0.75$, which leads to a QCP at $\kappa_{c}\approx 0.15$ marked by the vertical dashed gray lines in the following panels. (b) The QFI dynamics for different values of $\kappa$. The apparently enhanced QFI (dark blue curve) corresponds to the quench parameter $\kappa$ close to the QCP. (c) The time-averaged QFI exhibits a maximum in the vicinity of the QCP. (d) Histogram of $m=10000$ simulated measurement outcomes of $\hat{S}_{x}$ on the posted-quenched state at $t=7.5$. (e) The time average of quantum fluctuations in Eq. (6) is obtained for different measurement resolutions. They all show a peak that coincides well with the equilibrium QCP. (f) Histogram of 2000 simulated MLE of the interferometric phase using the posted-quenched state at $t=7.5$. The solid lines represent the Gaussian fitting centered around the true value $\vartheta=0.01$. (g) Time-averaged uncertainty of MLE estimate, see Eq. (8), identifies the QCP by its minimum. Here, the system size is set as $N=20$ and the window of the time-averaging operation is chosen as $T=15$.
Probing QPTs by global quantum fluctuations
The quantum fluctuations of the generator, defined as $\operatorname{Var}[\hat{\mathcal{O}}]=\langle\Psi|\hat{\mathcal{O}}^{2}|\Psi\rangle-\langle\Psi|\hat{\mathcal{O}}|\Psi\rangle^{2}$, are proportional to the QFI [see Eq. (1)]. It thus reflects the metrological capability of the post-quenched state. Therefore, a direct probing of QPTs via the metrologically capability can be implemented by the measurement of $\hat{\mathcal{O}}$ (e.g. $\hat{S}_{x}$ in ANNNI example) with respect to the post-quenched state |$\ket{\Psi_{t}}$. For a measurement repetitions of $m$ times, one obtains a series of random outcomes $\{S_{x}^{(1)},S_{x}^{(2)},\ldots,S_{x}^{(m)}\}$ taking (half-)integer values in $[-N/2,N/2]$. The fluctuations of $\hat{S}_{x}$ can be efficiently estimated using the Bessel Formula,
$\operatorname{Var}[\hat{S}_{x}]=\frac{1}{m-1}\sum_{i=1}^{m}\left(S_{x}^{(i)}-\langle\hat{S}_{x}\rangle\right)^{2},$ (6)
where $\langle\hat{S}_{x}\rangle=\sum_{i=1}^{m}S_{x}^{(i)}/m$. As demonstrated in Fig. 2(d), the measurement outcomes near the QCP apparently satisfy a wider probability distribution than the non-critical case. We calculate the time-averaged fluctuations $\overline{\operatorname{Var}\hat{S}_{x}}=(1/T)\int_{0}^{T}\operatorname{Var}[\hat{S}_{x}]\,\mathrm{d}t$, which evidently shows a peak versus the quench parameter $\kappa$ around the QCP, see Fig. 2(e).
To account for the influence of a limited experimental measurement resolution that technically coarsens the measurement outcomes, the outcome $S_{x}^{(i)}$ for the $i$-th measurement can be assumed to be continuously distributed, with the corresponding measurement operator given by [74, 75, 55, 73]
$\hat{\Pi}(S_{x}^{(i)})=\frac{1}{\sqrt{2\pi}\Delta}e^{-(S_{x}^{(i)}-\hat{S}_{x})^{2}/(2\Delta^{2})}.$ (7)
The standard deviation $\Delta$ effectively quantifies the measurement resolution, and when $\Delta\to 0$, Eq. (7) recovers the ideal projective measurement $\hat{\Pi}(S_{x}^{(i)})=\delta(S_{x}^{(i)}-\hat{S}_{x})$ with $\delta(x)$ denoting the Dirac delta function. We numerically simulate the measurement process of $\hat{S}_{x}$ under different resolutions and then evaluate quantum fluctuation based on Eq. (6). In Fig. 2 (e), we depict the time-averaged $\overline{\operatorname{Var}\hat{S}_{x}}$ versus the quench parameter $\kappa$. Our results shows that the position of the peak at QCP is not influenced even for a poor measurement resolution of $\Delta\sim\sqrt{N}\approx 4.5$, which would facilitate the observable $\operatorname{Var}[\hat{S}_{x}]$ to function as a trustworthy probe of QPTs.
Probing QPTs by estimation uncertainty
Naturally, the quantum probe state with a larger metrological capability gives to a better precision in estimating the phase $\vartheta$ associated with the unitary parameterization, i.e., $|\Psi_{t}(\vartheta)\rangle=e^{-i\vartheta\hat{\mathcal{O}}}|\Psi_{t}\rangle$. Stated by the so-called quantum Cramér-Rao bound, the best achievable estimation precision by detecting $n$ copies of $|\Psi_{t}(\vartheta)\rangle$ is lower bounded by $(\Delta\vartheta)^{2}\geq 1/[nF_{Q}(|\Psi_{t}\rangle,\hat{\mathcal{O}})]$. It should be pointed out that implementing the optimal measurement strategy to saturate such a bound is usually challenging. Here, we consider the experimentally feasible observables and show that the corresponding non-optimal detection in estimating $\vartheta$ can also witness the QPT.
For a general measurement, the outcomes obeys a $\vartheta$-dependent probability distribution, which determines the maximal information that can be extracted about $\vartheta$, i.e. $(\Delta\vartheta)^{2}\geq 1/(nF_{C})$, where $F_{C}$ denotes the Fisher information (FI) and $F_{C}\leq F_{Q}[|\Psi_{t}\rangle,\hat{\mathcal{O}}]$. This classical Cramér-Rao bound can be achieved by the maximum likelihood estimation (MLE) for a sufficiently large $n$ [76, 77, 73]. In the explicit example of ANNNI model, we estimate $\vartheta$ by suitably choosing to measure the global magnetization $\hat{S}_{z}=\sum_{i}\sigma_{i}^{z}/2$ [73]. We numerically simulate the estimation uncertainty of $\vartheta$ by utilizing the MLE, which is constructed from 50000 measurement outcomes. In order to evaluate the estimation uncertainty $\Delta\vartheta$, namely the fluctuation of the unbiased MLE, we repeat each estimation by 2000 times, from which the distribution of the MLE is obtained in Fig. 2 (f). It can be seen that the MLE near the QPT (corresponding to $\kappa\approx 0.15$) shows a narrower distribution than non-critical cases, directly indicating a better metrological performance achieved by the probe state generated at QCP. In Fig. 2 (g), we further compute the time average of the estimation uncertainty, i.e.
$\overline{\Delta\vartheta}=\frac{1}{T}\int_{0}^{T}\Delta\vartheta\,\mathrm{d}t.$ (8)
It clearly exhibits a minimum around the QCP. Hence, the parameter estimation uncertainty achieved by the probe state prepared through non-equilibrium quench dynamics effectively identifies the QPT.
Figure 3: Identification of the phase diagram of the ANNNI model. In the $(\kappa,B)$ parameter plane of the ANNNI chain, see Eq. (3), we numerically determine the boundary of the phase diagram by finding the maximum of the time-averaged QFI and the global quantum fluctuation [$\Delta=4.5$ in Eq. (7)], as well as the minimum of the time-averaged phase estimation uncertainty, see Fig.2. The boundaries obtained based on our dynamical quench-interferometric framework all agree well with the theoretical prediction by the second-order perturbation theory. Here, we set $N=20$. The $\overline{\operatorname{Var}(\hat{S}_{x})}$ and $\overline{\Delta\vartheta}$ are averaged over a finite time window $T=30$ with a sampling interval $\Delta t=0.3$ [73].
The feasible detection of quantum phase diagram
The quench-interferometric framework in Fig. 1 establishes a practical routine to identify QPTs. Based on this framework, we next elaborate that the quantum phase diagram can be obtained by sweeping the quenching parameter $(\kappa,B)$ in the ANNNI Hamiltonian. More specifically, we numerically determines $\kappa_{c}$ for each fixed $B$ and then varies $B$ from 0 to 1. In Fig. 3, we depict the phase diagram obtained from the metrological capability quantified by the QFI, the global quantum fluctuations and the phase estimation uncertainty. One can see that the predicted boundary all agrees very well with the theoretical prediction by second-order perturbation theory.
Recently, state-of-the-art quantum simulators have been implemented in a variety of well-controllable many-particle quantum platforms, such as the cold-atom lattice [78, 79, 80, 81], the Rydberg-atom array [82, 83, 84], and the trapped ion crystal [85, 86], which possess long coherence times, tunable interactions and starts to play significant roles in the simulation of a series of interesting spin models. As an example, we point out that the ANNNI chain considered in this work can be effectively synthesized using Rydberg atoms [23, 87, 88, 89, 90]. Particularly, the fully polarized initial state can be realized with a high fidelity of $\gtrsim 95\%$ [88, 89, 90]. The variance and the uncertainty of the MLE estimate can be readily obtained by simple global measurement of the collective spin operators [55, 91, 92, 93, 94, 95, 96, 97, 98]. Furthermore, the coherent timescale of this system can reach $\tau\geq 10J^{-1}$, where $J$ represents the strength of the nearest-neighbor interaction. Such a long relevant time scale makes the evaluation of the averaged metrological capability over a time window in our probing framework experimentally feasible. In the Supplementary Materials, we provide a detailed discussion on the operational complexity of the time-averaging requirement [73].
Conclusion &\&& outlook
To summarize, we present a quench-interferometric framework to probe QPTs by introducing the concept of the metrological capability to characterize the post-quenched state, which exhibit a significant peak around the QCP as quantified by the time average of QFI evolution. This framework can be accomplished through globally measuring collective observables to evaluate the quantum fluctuations or the associated estimation uncertainty of the interferometric phase. By taking the chaotic ANNNI model as a representative example, we demonstrate that the QCPs and immediately the phase diagram can be well identified based on such a framework. The scheme is experimentally feasible even with limited measurement resolution of the observables and under the implementation of non-optimal measurement to extract the phase information encoded in the final state. We also demonstrate that our framework can be extended to other QPT Hamiltonians including long-range interacting models as well [73]. Our findings take a step of exploiting quantum metrological concepts and strategies towards the investigation of many-body physics, which heuristically unveils a stimulating interplay between quantum metrology and dynamical quantum criticality [47, 39, 50].
Acknowledgements.
Acknowledgements
The computation is completed in the HPC Platform of Huazhong University of Science and Technology. This work is supported by the National Natural Science Foundation of China (12161141011 and 12174138), the National Key R&\&&D Program of China (2018YFA0306600). Y.-M.C. is also supported by the Young Scientists Fund of the National Natural Science Foundation of China (Grant No. 12304572).